OpenStudy (anonymous):

conjugate factor theorem.. does it apply for real roots? if 3+0i is a root 3-0i also a root ??

OpenStudy (anonymous):

but a polynomial only has n roots?? what if degree 2.. 3+0i and 3-0i both root how can u have the other or 3+0i and 3-0i count as one

OpenStudy (anonymous):

that is ur misconception if u find the root say (x-3)^2 it is having mathematically two roots 3,3

OpenStudy (anonymous):

whut .. thats caue multiciplity =2 okay take x^2+4x+3=0 (x-3)(x-1)=0 x=3,1

OpenStudy (anonymous):

no u have not taken my point in depth see if u shift this parabola y=(x-3)^2 by 1 unit down u get two roots but if it was having single root then it will never be the case that matters a lot!!

OpenStudy (anonymous):

huh o-o yeah so for my example.. youre saying that the roots would be x=3,3,1,1 because they have to occur in conjugagtes ?

OpenStudy (anonymous):

no it is not the case in ur second reply but it was indeed in first reply because there conjugate was hidden in the quadratic formula(in ur second reply)

OpenStudy (anonymous):

what do u mean hidden ??

OpenStudy (anonymous):

i mean when we write quadratic formula for ax^2 +bx +c=0 then (-b(+/-)(b^2-4ac)^1/2)/2a the term in square root is the conjugate term

OpenStudy (anonymous):

it can have conjugate by taking for one time + and another time -

OpenStudy (anonymous):

r u getting it or not

hartnn (hartnn):

in short, conjugate theorem does not apply to real roots, if 3+0i is factor, u cannot say 3-0i is also the factor......what else is the confusion?

OpenStudy (anonymous):

no hartnn mathematically they are two different roots with same MAGNITUDE

hartnn (hartnn):

nopes, they are same roots.....i mean 3+0i=3-0i=3 and u cannot consider magnitude or angle as both are real numbers not complex at all.

hartnn (hartnn):

the magnitude of a real number does not quite make sense if defined,its value is same a that of the number.

OpenStudy (anonymous):

i do not believe if shift the graph of y=(x-3)^2 one unit down then it has two roots which can not be the case if it has only distinct roots any such graphical trans formation can not change the number of roots

hartnn (hartnn):

if u shift the graph of y=(x-3)^2 one unit down then the new equation will be y-1=(x-3)^2 as all the points now have their new y co-ordinate as y-1 \((x-3)^2=0\) had two equal solutions 3,3 -?grapg cutting at only x=3 \((x-3)^2+1=0\) will have 2 distinct solutions which implies graph cuts at 2 distinct points............

hartnn (hartnn):

when u are shifting the graph,the equation of graph changes and so does the intercepts , got it @RajshikharGupta ?

OpenStudy (anonymous):

alright thanks hartnn :)

OpenStudy (anonymous):

isnt equation y=(x-3)^2-1 if u shift down ?

hartnn (hartnn):

oh yes, thanks for pointing it out Y=y-1 so y=Y+1 so new equation: Y+1=(X-3)^2 or Y=(X-3)^2+1

hartnn (hartnn):

sorry Y=(X-3)^2-1

OpenStudy (anonymous):

i m not satisfied with ur explanation out of my intuitions!!!

hartnn (hartnn):

so just plot it and see... y=(x-3)^2 and y=(x-3)^2-1

OpenStudy (anonymous):

 same graph different roots

hartnn (hartnn):

ofcourse same graph will have different roots(roots= x intercepts) depending on where u place the graph.......